# A Comprehensive Survey of Artificial Intelligence Techniques for Talent Analytics

### v2

@@ Abstract
In today’s competitive and fast-evolving business environment, it is a critical time for organizations to rethink how to make talent-related decisions in a quantitative manner. Indeed, the recent development of Big Data and Artificial Intelligence (AI) techniques have revolutionized human resource management. The availability of large-scale talent and management-related data provides unparalleled opportunities for business leaders to comprehend organizational behaviors and gain tangible knowledge from a data science perspective, which in turn delivers intelligence for real-time decision-making and effective talent management at work for their organizations. In the last decade, talent analytics has emerged as a promising field in applied data science for human resource management, garnering significant attention from AI communities and inspiring numerous research efforts. To this end, we present an up-to-date and comprehensive survey on AI technologies used for talent analytics in the field of human resource management. Specifically, we first provide the background knowledge of talent analytics and categorize various pertinent data. Subsequently, we offer a comprehensive taxonomy of relevant research efforts, categorized based on three distinct application-driven scenarios from different level: talent management, organization management, and labor market analysis. In conclusion, we summarize the open challenges and potential prospects for future research directions in the domain of AI-driven talent analytics.
@@ 1 Introduction
In the world of volatility, uncertainty, complexity, and ambiguity (VUCA), talents are always precious treasures and play an important role for business success. To cope with the fast-evolving business environment and maintain competitive edges, it is critical for organizations to rethink how to make talent-related decisions in a quantitative manner. Thanks to the era of big data, the availability of large-scale talent data provides unparalleled opportunities for business leaders to understand the rules of talent and management, which in turn deliver intelligence for effective decision-making and management for their organizations [374, 361]. Along this line, as an emerging applied data science direction in human resource management, talent analytics has attracted a wide range of attention from both academic and industry circles. Specifically, talent analytics, also as known as workforce analysis or people analytics, focuses on leveraging data science technologies to analyze extensive sets of talent-related data, empowering organizations with informed decision-making capabilities that enhance their organizational and operational effectiveness. [206]. In practice, talent analytics plays a pivotal role in strategic human resource management (HRM), encompassing diverse applications such as talent acquisition, development, retention, as well as examining organizational behaviors and external labor market dynamics. Generally, the research directions of talent analytics can be divided into three categories, as illustrated in Figure 1, including talent management, organization management, and labor market analysis.
@@
To be specific, first, talent management is a constant strategic process of attracting and hiring the high-potential employees, training their skills, motivating them to improve their performance, and retaining them to keep organizational competitiveness. In this particular scenario, talent analytics primarily focus on individual-level analysis. For instance, it can help human resources managers find the right talents for different jobs in a practical way [334, 140, 444], and it can reasonably make the employee performance or turnover prediction [6, 233]. Second, organization management is the art of fostering collaboration among talents and guiding the organizations toward achieving success. In this scenario, talent analytics can diagnose the health of an organization and measure the organizational performance by leveraging various relationship information between talents or organizations, such as organizational structure, communication patterns, and project collaborations [446, 114]. It can also assist the organization in effectively structuring and optimizing teams [414, 13]. Third, talent analytics can be applied from an external and macro perspective, i.e., applied to labor market analysis scenario. It is crucial to devise talent and organizational strategies. For instance, by analyzing talent demands within the labor market, managers can effectively craft recruitment strategies [484, 467].
@@
Historically, talent analytics is proposed within the conception of HRM around 1920s [191]. Talent analytics is usually manual in the early stages of HRM before 1970 [280]. In this stage, human resource systems (HRS) or human resource management systems (HRMS) is the main method for talent analytics [51]. There are various types of these systems like commitment system or control system [25]. The assessment of the talent performance with these systems are conducted by the human resource managers [138]. At the same time, talent assessment-related theories have formed like human capital theory and resource-based theory [191]. With the occurrence of mechanical automation, human resource information systems (HRIS) has arised around 1940s, however, before 1970, HRIS is based on the sorting and tabulating equipment, at the same time, the main functions of HRIS are keeping employee records automatically and there is no computer support [111]. However, from this stage, talent analytics could be supported by aggregated information. With the development of information systems motivated by more and more data generated during management, HRIS enhanced by computer system have been widely adopted in 1970s [59], in this stage, HRIS is a combination of the database, the computer application, the software and the hardware to record, manage, and operate human resource data. This development trend is verified by a survey, which confirms that 60% of Fortune 500 companies use HRIS to support daily HRM operations [36]. Besides, in this stage, people analytics, refers to a novel, quantitative, evidence-based, and data-driven approach to manage the workforce, is proposed to raise the efficiency of core human resource functions such as recruiting [142]. Some standard statistical methods are adopted in this step like correlation analysis, simple regressions and so on [229], this way is generally called descriptive analytics. At the same time, due to the rich functions of HRIS, many evaluation studies on the specific functions of software have been carried out, including organizational performance, turnover and so on [126, 21].
@@
With the development of artificial intelligence (AI) algorithms, some advanced regression techniques, data mining, text mining, web mining, and forecast calculations are used in talent analytics around 2010s [117], these ways is generally called predictive analytics. Recent literature highlights the importance of more advanced analytical methods and emerging technologies in talent analytics. For instance, Dahlbom et al. [94] emphasise that “new types of data and different algorithms used in AI and machine learning solutions utilized in HRA [Human Resource Analytics]” (p. 123) will transform the field of people analytics. There are two main challenges to facilitate this transformation. On the one hand, talent analytics is facing digital disruption, which enables the availability of large-scale relevant data. For instance, Indeed, a world-renowned job search site, had 11.3 million active jobs as of January 2022 [61]. Meanwhile, Linkedin, the largest online professional network, had 774 million members from around 200 countries as of March 2022 [248], building up a wealth of labor market data. Moreover, numerous enterprises are setting up their Digital Human Resource Management Systems (Digital HRMS), enabling the collection, storage, and processing of a huge amount of talent and organizational information in a digital environment [393]. On the other hand, with the advent of talent-related big data, advanced AI techniques have rapidly revolutionized a series of research and practices in this field at an alarming rate, which in turn deliver intelligence for decision-making and management for their organizations. In this stage, the deep learning methods have enabled the new paradigm in person-job fit [334, 485, 46, 444] and person-organization fit [381, 383], so as to achieve the efficient and accurate talent selection and development. Text mining methods have been adopted in the employer brand analysis based on the large-scale labor market data [246, 245], which enable the forward-looking strategic plans created for the business. At the same time, several high-tech companies are gradually incorporating AI technologies into their HRMS. As an illustration, IBM leverages AI technology to achieve a remarkable 95 percent accuracy in predicting employees who are considering leaving their positions, which saved IBM \$300 million in retention costs [345]. With the strong automation capability of large language model (LLM), autonomous people analytics is conducted around 2020s [315].
@@
Accordingly, AI in talent analytics in this paper includes supervised learning, unsupervised learning, deep learning, reinforcement learning, knowledge representation, natural language processing, and so on [43]. These techniques construct different business capabilities in people analytics, which are automation of structured (or semistructured) work processes, engagement with employees and managers, decision-making through extensive analysis of a large amount of data, creation of novel outcomes [43]. This survey attempts to provide a comprehensive review of the rapidly evolving AI techniques for talent analytics. Based on our investigation, we first provide a detailed taxonomy of relevant data laying a data foundation for leveraging AI techniques to understand talents, organizations, and corresponding management better. Generally, talents’ behaviors reflect in three levels, including individual level, organizational level and market level. Accordingly, the research efforts of the AI techniques for talent analytics from corresponding three aspects, including talent management, organization management, and labor market analysis. Finally, we identify challenges for future AI-based talent analytics and suggest potential research directions.
Moreover, in order to help the readers learn more effectively, we highlight the systematic resources provided in this survey as follows,
  • Table I summarizes the data for talent analytics.
  • Table III summarizes the recent AI-based talent analytics efforts in the talent management scenario.
  • Table IV summarizes the recent AI-based talent analytics efforts in the organization management scenario.
  • Table V summarizes the recent AI-based talent analytics efforts in the labor market analysis scenario.
@@ 2 Data for Talent Analytics
Nowadays, as enterprises undergo an accelerated digital transformation, a large amount of talent analytics-related data has been accumulated. In this section, we will introduce the data collected across various scenarios, providing readers with a foundational understanding of the related research data and the motivation for model design. Generally, the data can be divided into internal data, which are collected from the internal enterprise management system, and external data, which are collected from the external labor market.
@@ 2.1.1 Recruitment Data
Resume: A resume or Curriculum Vitae (CV) is a document that outlines a person’s background, skills, and accomplishments, which plays a vital role in the recruitment process as it serves to facilitate talent screening and assessment [334, 363]. It serves as an important tool for job seekers to showcase their qualifications and suitability for a job position. Recently, a large amount of resume data, in either Word or PDF format, has accumulated with the development of online recruitment. As shown in Figure 3, a resume typically comprises structured information such as gender, age, and education, as well as semi-structured information like educational experience, work experience, and project experience. Accordingly, several resume parsing techniques have been developed to extract the redundant information  [74, 75, 443]. On such basis, substantial efforts are posed in talent analytics with resume data from different perspectives. For instance, Yao et al. [442] introduced a keyphrase extraction approach to explore job seekers’ skills in resumes, and Pena et al. [323] used image information in resume data to improve screening performance and explore fairness issue. Moreover, several studies have proposed leveraging the text mining techniques to determine the matching degree between jobs and job seekers based to their resumes [334, 485, 46]. In addition, the resumes also encompass the career trajectories of the job seekers. As illustrated on the right side of Figure 3, the candidate’s profile showcases three job experiences, including a job change at Microsoft and work experience at Google. In this phase, Zhang et al. introduced the ResumeVis system to visualize the individual career trajectory and mobility within different organizations [460]. And the researchers further analyzed the sequential patterns of the career trajectory and proposed personalized career development recommendation [289, 465, 400].
@@
Job Posting: A job posting is an advertisement for a vacant position that provides job seekers with information on the job description and requirements. The posting offers applicants a clear understanding of what the position is responsible for and what qualifications are necessary. Recently, the proliferation of online recruitment services has made it increasingly common to publish job postings as web pages. Figure 3 illustrates a typical job posting that comprises structured information, such as the salary range and education requirements, as well as the semi-structured content that includes job duties descriptions and abilities requirements. Nevertheless, it is still difficult to deal with such a large corpus of data by Human Resource (HR) experts manually. To this end, researchers have attempted to reduce the dependence on manual labor by using neural network-based techniques, particularly NLP, on voluminous job postings. As mentioned before, considerable effort has been posed in Person-Job Fit  [334, 335, 46, 485, 263], which aims to match the job postings with suitable resumes. Moreover, Shen et al. leveraged the latent variable model to jointly model the job description, candidate resume, and interview assessment, which can further benefit several downstream applications such as person-job fit and interview question recommendation [364, 363]. In order to reduce the expense of manual screening, researchers also extracted the job entities from postings and generated interview questions automatically  [336, 365, 333, 239]. Apart from these in-firm applications, some studies are carried out to provide comprehensive insights into the global labor market. For instance, researchers have proposed several data-driven methods for salary analysis across different companies and positions [210, 82, 288]. Zhang et al. utilized large-scale job postings from one of the largest Chinese online recruitment websites and forecast fine-grained talent demand in the recruitment market [467]. Moreover, some studies aim to measure the popularity of job skills and forecast their evolving trends [426, 422]. Along this line, Sun et al. further focus on measuring the values of job skills based on massive job postings, contributing to the quantitative assessment of job skills [382].
@@
Interview-related Data: Interview-related data is typically collected during the interview process and serves the purpose of evaluating applicants’ overall qualifications for the position they are applying for. In general, an interview can be conducted either in-person or through video, resulting in textual or video-based assessments, respectively. Both of these two kinds of data enable comprehensive evaluations for the candidates and facilitate the integration of AI in HR. To address the subjectivity of traditional interviews, Shen et al. utilized the latent variable model to explore the relationship among job descriptions, candidate resumes, and textual interview assessments [364, 363]. The results provide an interpretable understanding of job interview assessments. Indeed, textual interview assessments within a company are usually private and sensitive, whereas video assessments draw more attention. For instance, several studies extract multimodal features from the videos for automatic analysis of job interviews [78, 80]. In addition, Hemamou et al. proposed a hierarchical attention model to predict the employability of the candidates using multimodal information, including text, audio, and video [172]. Along this line, Chen et al. leveraged a hierarchical reasoning graph neural network to automatically score candidate competencies using textual features in asynchronous video interviews [77].
@@ 2.1.2 Employee Data
Regarding the development of employees within a company, a significant amount of employee data has been accumulated, including training records and individual work outcomes. An overview of employee-related data is shown in Figure 4.
Employee Profiles: Employee profiles typically describe an individual based on two main aspects: demographic characteristics and individual work outcomes. In specific, the former branch includes characteristics such as age, gender, and education levels [322], which can be used to enhance employee representations and benefit various downstream analyses, such as career mobility prediction  [337, 477] and performance forecasting [271]. In addition to these static variables, individual work outcomes depict the dynamic career development from different dimensions, such as performance appraisals, promotion, and turnover records. In particular, a performance appraisal is a systematic evaluation of an employee’s job performance and productivity that is typically conducted by line managers. Besides, the promotion and turnover records show employee movements within and across companies, respectively. All of this information contributes to further insights into employee dynamics. For instance, researchers have leveraged performance appraisals to identify the high-potential talents within a company [447]. Li et al. utilized static profiles, performance appraisals, and reporting lines of employees to model career development within a company, focusing on turnover and career progression [233]. Sun et al. proposed to capture the dynamic nature of person-organization fit based on individual profiles, reporting lines, and communication records [381]. To investigate the contagious effect of turnovers, researchers have utilized both employee profiles and turnover data [387, 386]. Furthermore, Hang et al. leveraged five kinds of standardized data, including employee turnover records, to predict the turnover probability and period [165].
@@
Training Records: Employee training is a program designed to improve the performance of employees by equipping them with specific skills. Ongoing employee training has proven to be crucial in attracting and retaining top talent [351]. Typically, the training record describes the learning path of an employee, which is a sequence of different skills. Based on these training records, considerable effort has been devoted to exploring the learning patterns of employees. For instance, Wang et al. utilized both learning records and skill profiles of employees from a high-tech company in China to develop a personalized online course recommendation system [402, 401]. Along this line, Srivastava et al. collected employees’ training and work history from a large multinational IT organization to provide personalized next training recommendations [377]. In addition, some researchers also provided insights into employee competency study [234, 421, 252, 177]. For instance, multi-dimensional features, including learning and training dimensions, were collected from a Chinese state-owned enterprise to provide competency assessment for employees [252].

### v3

@@ Abstract
In today’s competitive and fast-evolving business environment, it is critical for organizations to rethink how to make talent-related decisions in a quantitative manner. Indeed, the recent development of Big Data and Artificial Intelligence (AI) techniques has revolutionized human resource management. The availability of large-scale talent and management-related data provides unparalleled opportunities for business leaders to comprehend organizational behaviors and gain tangible knowledge from a data science perspective, which in turn delivers intelligence for real-time decision-making and effective talent management for their organizations. In the last decade, talent analytics has emerged as a promising field in applied data science for human resource management, garnering significant attention from AI communities and inspiring numerous research efforts. To this end, we present an up-to-date and comprehensive survey on AI technologies used for talent analytics in the field of human resource management. Specifically, we first provide the background knowledge of talent analytics and categorize various pertinent data. Subsequently, we offer a comprehensive taxonomy of relevant research efforts, categorized based on three distinct application-driven scenarios at different levels: talent management, organization management, and labor market analysis. In conclusion, we summarize the open challenges and potential prospects for future research directions in the domain of AI-driven talent analytics.
@@ 1 Introduction
In the world of volatility, uncertainty, complexity, and ambiguity (VUCA), talent are always precious treasures and play an important role in business success. To cope with the fast-evolving business environment and maintain competitive edges, it is critical for organizations to rethink how to make talent-related decisions in a quantitative manner. Thanks to the era of big data, the availability of large-scale talent data provides unparalleled opportunities for business leaders to understand the patterns of talent and management, which in turn deliver intelligence for effective decision-making and management for their organizations [1, 2]. Along this line, as an emerging applied data science direction in human resource management, talent analytics has attracted a wide range of attention from both academic and industry circles. Specifically, talent analytics, also known as workforce analysis or people analytics, focuses on leveraging data science technologies to analyze extensive sets of talent-related data, empowering organizations with informed decision-making capabilities that enhance their organizational and operational effectiveness [3]. In practice, talent analytics plays a pivotal role in strategic human resource management (HRM), encompassing diverse applications such as talent acquisition, development, retention, as well as examining organizational behaviors and external labor market dynamics. Generally, the research directions of talent analytics can be divided into three categories, as illustrated in Figure 1, including talent management, organization management, and labor market analysis.
@@
To be specific, first, talent management is a constant strategic process of attracting and hiring high-potential employees, training them in needed skills, motivating them to improve their performance, and retaining them to keep organizational competitiveness. In this particular scenario, talent analytics primarily focuses on individual-level analysis. For instance, it can help human resource managers find the right talents for different jobs in a practical way [4, 5], and it can reasonably make employee performance or turnover predictions [6]. Second, organization management is the art of fostering collaboration among talent and guiding organizations toward achieving success. In this scenario, talent analytics can diagnose the health of an organization and measure organizational performance by leveraging various relationship information among talents and organizations, such as organizational structure, communication patterns, and project collaborations [7]. It can also assist the organization in effectively structuring and optimizing teams [8]. Third, talent analytics can be leveraged from an external, macro-level perspective, particularly in labor market analysis scenarios. It is crucial to devise talent and organizational strategies. For instance, by analyzing talent demands within the labor market, managers can effectively craft recruitment strategies [9, 10].
@@
Historically, talent analytics was proposed within the conception of HRM around the 1920s [11]. Talent analytics is usually manual in the early stages of HRM before 1970 [12]. At this stage, Human Resource Systems (HRS) or Human Resource Management Systems (HRMS) serve as the primary tools for talent analytics. [13]. There are various types of these systems, such as commitment or control systems [14]. The assessment of the talent performance with these systems is conducted by the human resource managers [15]. At the same time, talent assessment-related theories have emerged, including human capital theory and resource-based theory [11]. With the occurrence of mechanical automation, human resource information systems (HRIS) arose around the 1940s; however, before 1970, HRIS was based on sorting and tabulating equipment. At the same time, the main functions of HRIS are keeping employee records automatically, and there is no computer support [16]. However, from this stage, talent analytics could be supported by aggregated information. With the development of information systems motivated by more and more data generated during management, HRIS enhanced by computer system has been widely adopted in 1970s [17], in this stage, HRIS is a combination of the database, the computer application, the software and the hardware to record, manage, and operate human resource data. This development trend is verified by a survey, which confirms that 60% of Fortune 500 companies use HRIS to support daily HRM operations [18]. Besides, in this stage, people analytics, refer to a novel, quantitative, evidence-based, and data-driven approach to manage the workforce, was proposed to raise the efficiency of core human resource functions such as recruiting [19]. Some standard statistical methods were adopted in this step, such as correlation analysis and simple regressions [20]; this is generally referred to as descriptive analytics. At the same time, due to the rich functions of HRIS, many evaluation studies on the specific functions of software have been carried out, such as organizational performance, turnover, and so on.
@@
With the development of artificial intelligence (AI) algorithms, some advanced regression techniques, data mining, text mining, web mining, and forecast calculations have been used in talent analytics around the 2010s. On the one hand, talent analytics is facing digital disruption, which enables the availability of large-scale relevant data. For instance, a world-renowned job search site had 11.3 million active jobs as of January 2022 [21]. Meanwhile, LinkedIn, the largest online professional network, had 774 million members from around 200 countries as of March 2022 [22], building up a wealth of labor market data. Moreover, numerous enterprises are setting up their Digital Human Resource Management Systems (Digital HRMS), enabling the collection, storage, and processing of a huge amount of talent and organizational information in a digital environment [23]. On the other hand, with the advent of talent-related big data, advanced AI techniques have rapidly revolutionized a series of practices in this field at an alarming rate, which in turn deliver intelligence for decision-making and management for their organizations. In this stage, deep learning methods have enabled the new paradigm in person-job fit [4, 24] and person-organization fit [25, 26], so as to achieve the efficient and accurate talent selection and development. Text mining methods have been adopted in the employer brand analysis based on the large-scale labor market data [27, 28], which enables the forward-looking strategic plans created for the business. At the same time, several high-tech companies are gradually incorporating AI technologies into their HRMS. With the strong automation capability of large language models (LLMs), autonomous people analytics was introduced around the 2020s [29].
@@
Accordingly, AI in talent analytics in this paper includes supervised learning, unsupervised learning, deep learning, reinforcement learning, knowledge representation, natural language processing, and so on [30]. These techniques construct different business capabilities in people analytics, which are automation of structured (or semistructured) work processes, engagement with employees and managers, decision-making through extensive analysis of a large amount of data, and creation of novel outcomes [30]. This survey attempts to provide a comprehensive review of the rapidly evolving AI techniques for talent analytics. Based on our investigation, we first provide a detailed taxonomy of relevant data, which lays a foundation for better leveraging AI techniques to understand talent, organizations, and their management. Generally, talent behaviors are reflected in three levels, including the individual level, organizational level, and market level. Accordingly, the research efforts of the AI techniques for talent analytics from the corresponding three aspects, including talent management, organization management, and labor market analysis. Finally, we identify challenges for future AI-based talent analytics and suggest potential research directions.
@@ 2 Data for Talent Analytics
As enterprises undergo accelerated digital transformation, a large amount of data related to talent analytics has been accumulated. In this section, we introduce the data collected from various scenarios to provide readers with a foundational understanding of the research data and the motivations behind model design. Generally, the data can be classified into internal data (collected from enterprise management systems) and external data (collected from the external labor market).
@@ 2.1.1 Recruitment Data
Resume: A resume or Curriculum Vitae (CV) is a document that outlines a person’s background, skills, and accomplishments, which plays a vital role in the recruitment process as it serves to facilitate talent screening and assessment [4, 31]. It is an important tool for job seekers to demonstrate their qualifications and fit for the position. Driven by the development of online recruitment, a large amount of resume data, typically in Word or PDF format, has recently been accumulated. As shown in Figure 3, a resume typically comprises structured information such as gender, age, and education, as well as semi-structured information like educational experience, work experience, and project experience. Accordingly, several resume parsing techniques have been developed to extract the redundant information  [32, 33]. On this basis, substantial efforts have been made in talent analytics with resume data from different perspectives. For instance, Yao et al. [34] introduced a keyphrase extraction approach to explore job seekers’ skills in resumes, and Pena et al. [35] used image information in resume data to improve screening performance and explore the fairness issue. Moreover, several studies have proposed leveraging the text mining techniques to determine the matching degree between jobs and job seekers based on their resumes [4, 36]. In addition, the resumes also encompass the career trajectories of the job seekers. As illustrated on the right side of Figure 3, the candidate’s profile showcases three job experiences, including a job change at Microsoft and work experience at Google. In this phase, Zhang et al. introduced the ResumeVis system to visualize the individual career trajectory and mobility within different organizations [37]. And the researchers further analyzed the sequential patterns of the career trajectory and proposed personalized career development recommendations [38].
The presence of substantial sensitive personal information in resume data poses significant challenges to the creation of publicly accessible datasets. While some researchers have used publicly available information from sources such as academic homepages and online professional networks to build alternative datasets [39, 40], the severe lack of open datasets continues to hinder broader research in this area.
@@
Job Posting: A job posting is an advertisement for a vacant position that provides job seekers with information on the job description and requirements. The posting offers applicants a clear understanding of what the position is responsible for and what qualifications are necessary. Recently, the proliferation of online recruitment services has made it increasingly common for job postings to be published as web pages. Figure 3 illustrates a typical job posting that comprises structured information, such as the salary range and education requirements, as well as semi-structured content that includes job duty descriptions and ability requirements. Nevertheless, it is still difficult to deal with such a large corpus of data for Human Resource (HR) experts manually. To this end, researchers have attempted to reduce the dependence on manual labor by using neural network-based techniques, particularly NLP, on voluminous job postings. As mentioned before, considerable effort has been made in Person-Job Fit  [4, 41, 24], which aims to match the job postings with suitable resumes. Moreover, Shen et al. leveraged the latent variable model to jointly model the job description, candidate resume, and interview assessment, which can further benefit several downstream applications such as person-job fit and interview question recommendation [31]. To reduce the expense of manual screening, researchers also extracted the job entities from postings and generated interview questions automatically [42, 43, 44]. Apart from these in-firm applications, some studies are carried out to provide comprehensive insights into the global labor market. For instance, researchers have proposed several data-driven methods for salary analysis across different companies and positions [45, 46]. Zhang et al. utilized large-scale job postings from one of the largest Chinese online recruitment websites and forecast fine-grained talent demand in the recruitment market [10]. Moreover, several studies seek to measure the popularity of job skills and predict their evolving trends over time [47]. Along this line, Sun et al. further focus on measuring the values of job skills based on massive job postings, contributing to the quantitative assessment of job skills [48].
@@
Interview-related Data: Data collected during interviews, whether conducted in person or via video, plays a critical role in evaluating applicants’ overall qualifications for their intended positions. Textual and video-based assessments derived from these processes enable comprehensive candidate evaluations and promote the integration of AI into HR management practices. To address the subjectivity of traditional interviews, Shen et al. utilized the latent variable model to explore the relationship among job descriptions, candidate resumes, and textual interview assessments [49, 31]. The results provide an interpretable understanding of job interview assessments. Indeed, textual interview assessments within a company are usually private and sensitive, whereas video assessments draw more attention. For instance, several studies extract multimodal features from the videos for automatic analysis of job interviews [50]. In addition, Hemamou et al. proposed a hierarchical attention model to predict the employability of the candidates using multimodal information, including text, audio, and video [51]. Along this line, Chen et al. leveraged a hierarchical reasoning graph neural network to automatically score candidate competencies using textual features in asynchronous video interviews [52]. It is important to note that the use of interview data must account for potential ethical risks. Although researchers typically state that they have obtained permission to use data in their papers, usage of interview data should be approached with caution in real-world scenarios.
@@ 2.1.2 Employee Data
Regarding the development of employees within a company, a significant amount of employee data has been accumulated, such as training records and individual work outcomes, as illustrated in Figure 4.
Employee Profiles: Employee profiles typically describe an individual based on two main aspects: demographic characteristics and individual work outcomes. Specifically, the former branch includes characteristics such as age, gender, and education levels, which can be used to enhance employee representations and benefit various downstream analyses, such as career mobility prediction [53] and performance forecasting [54]. In addition to these static variables, individual work outcomes depict the dynamic career development from different dimensions, such as performance appraisals, promotion, and turnover records. In particular, a performance appraisal refers to the systematic evaluation of an employee’s job performance and productivity, typically conducted by line managers. In addition, promotion and turnover records reflect employee movements within and across companies, respectively. All of this information contributes to further insights into employee dynamics. For instance, researchers have leveraged performance appraisals to identify the high-potential talents within a company [55]. Li et al. utilized employees’ static profiles, performance appraisals, and reporting lines to model career development within a company, focusing on turnover and career progression [6]. Sun et al. proposed to capture the dynamic nature of person-organization fit based on individual profiles, reporting lines, and communication records [25]. To investigate the contagious effect of turnovers, researchers have utilized both employee profiles and turnover data [56]. Furthermore, Hang et al. leveraged five kinds of standardized data, including employee turnover records, to predict the turnover probability and period [57].
@@
Training Records: Employee training is a program designed to improve the performance of employees by equipping them with specific skills. Ongoing employee training has proven to be crucial in attracting and retaining top talent. Typically, the training record describes the learning path of an employee, which is a sequence of different skills. Based on these training records, considerable effort has been devoted to exploring the learning patterns of employees. For instance, Wang et al. utilized both learning records and skill profiles of employees from a high-tech company in China to develop a personalized online course recommendation system [58]. Along this line, Srivastava et al. collected employees’ training and work history from a large multinational IT organization to provide personalized next training recommendations [59]. Furthermore, several researchers have contributed insights into the study of employee competencies [60, 61]. For instance, multi-dimensional features, including learning and training dimensions, were collected from a Chinese state-owned enterprise to provide competency assessment for employees [61].
Due to legal regulations, employee profiles and training records, which encompass both employee privacy and the company’s proprietary training strategies, should not be made available. This limits the ability of the field to gain broader research attention. Furthermore, many existing studies on employee datasets lack discussion on gender and racial diversity, which hinders the ability to perform fairness assessments within these datasets. This also makes it challenging to identify and address potential biases that may arise in career mobility predictions, performance evaluations, or training outcomes for different demographic groups. As a result, research aimed at understanding the disparate impact of HR decisions across diverse employee populations remains constrained, hindering efforts to ensure equitable career development opportunities and unbiased workforce management strategies.